Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is “no solution” or “infinitely many solutions.”
5x – 4y = 1
-10x + 8y = -3

Answer by checkley77(12844)

5x–4y=1 multiply this equation by 2 & add them.
-10x+8y=-3
10x-8y=2
-----------------------
0x+0y=-1 indicates there is no solution.